this year real GDP per person in olympus is 45,000 while the year before it was 48,250. What is the growth rate of real GDP per Person
1
Expert's answer
2013-04-15T09:40:42-0400
Income per person, is the mean income
within an economic aggregate such as a country or city. It is calculated by
taking a measure of all sources of income in the aggregate (such as GDP or
Gross national income) and dividing it by the total population.
GDP per capita growth = (GDP2 - GDP1)/GDP1*100% = (45,000 - 48,250)/45,000*100%
= -7.22%,
So, GDP per capita decreased and the growth rate is negative.
this year real GDP per person in olympus is45,000 while the year before it
was 48,250. What is the growth rate of real GDP per Person